OpenStudy (anonymous):

A bit confused. A landscaper plans to dig a rectangular garden whose length is to be 5 feet more than the width. If the landscaper has 42 feet of fencing to enclose the garden, what should be the dimensions of the garden? I know I have to make 2 equations but not sure where to start ):

OpenStudy (anonymous):

length = width + 5 perimeter = 2(length + width )

OpenStudy (anonymous):

Okay thanks! Where do I go from now though...?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

put length = width + 5 in second equation

OpenStudy (anonymous):

Now Perimeter = 2(width +(width+5) ) Perimeter is given so solve for width

OpenStudy (anonymous):

It comes out being P=w+5?

OpenStudy (anonymous):

P = 2(w+w+5) P= 2(2w+5)

OpenStudy (anonymous):

Okay so P=4w+5

OpenStudy (anonymous):

And P=42 right?

OpenStudy (anonymous):

I got the W should be 9.25

OpenStudy (anonymous):

P= 2*2w +2*5 (distributive roperty) P = 42 42 = 4w +10 42-10= 4w so w= ??

OpenStudy (anonymous):

Oops I messed up one sec

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

So it's a 13x8 in diminsions?

OpenStudy (anonymous):

Correct. :)

OpenStudy (anonymous):

Awesome thanks so much!(:

OpenStudy (anonymous):

anytime:)

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

Sure. :)